# State of Health Estimation for Battery Modules with Parallel-Connected Cells Under Cell-to-Cell Variations

### 08 Dec 2023 v2

@@ Abstract
State of health (SOH) estimation for lithium-ion battery modules with cells connected in parallel is a challenging problem, especially with cell-to-cell variations. Incremental capacity analysis (ICA) and differential voltage analysis (DVA) are effective at the cell level, but they cannot be directly applied to module level SOH estimation, when only module level measurements are available. This paper proposes a new method and demonstrates that, with multiple features systematically selected from the module level ICA and DVA, the module level SOH can be estimated with high accuracy and confidence in the presence of cell-to-cell variations. First, a new information theory-based feature selection algorithm is proposed to find an optimal set of features for module level SOH estimation. Second, a new relevance vector regression (RVR)-based module level SOH estimation model is proposed to provide both point estimates and three-sigma credible intervals while maintaining model sparsity. Experimental datasets are used to illustrate and evaluate the proposed method. With more selected features incorporated, the proposed method achieves better estimation accuracy and higher confidence at the expense of higher model complexity. This trade-off is explored through a case study. When applied to a large experimental dataset, the proposed method and the resulting sparse model lead to module level SOH estimates with 0.5% root-mean-square errors and 1.5% average three-sigma values. With all the optimization and training processes completed offboard, the proposed method has low computational complexity for onboard implementations.
@@ I INTRODUCTION
Lithium-ion battery state of health (SOH) estimation with onboard measurements is an active research topic for electrified vehicles [1]. SOH, representing the degradation status of batteries, is critical information for battery management systems (BMS) [2]. Depending on the applications, SOH may be defined as resistance rising, capacity fading, or their combinations [3, 4]. This paper focuses on capacity fading. In this definition, degradation in SOH results in the range reduction of electrified vehicles [4].
While there is a rich set of methodologies and results in the literature for SOH estimation, most of the work reported focuses on the cell level [2, 3, 5], using model-based and data-driven methods. Popular model-based methods include Bayesian filters [6, 7, 8], sliding-mode observers [9], and least-square techniques [10]. However, given the complicated degradation process, robust and accurate estimation is difficult to achieve with simple phenomenological or physical models using only onboard measurements. On the other hand, data-driven methods have the advantage of not requiring physics-based models. Various reported data-driven methods include convolutional and recurrent neural networks [11, 12]. Without insights and structures from physics-based models, however, the effectiveness of these methods solely depends on the comprehensiveness and quality of training data.
@@
Figure 1: Example Module Level IC Curves around an IC Peak under Cell-to-Cell Variations
One promising middle ground between model-based and data-driven methods is incremental capacity analysis (ICA) [13, 14, 15]. Peaks in incremental capacity (IC) curves reflect degradation mechanisms inside batteries [16, 17]. Thus, ICA incorporates the physical features related to aging without using complicated mechanistic models. Furthermore, this method leverages data-driven techniques to correlate physical features (IC peaks) and SOH, and uses the learned model to estimate SOH [18]. ICA has been shown to be effective for different battery chemistries under different temperatures, C rates, and charging ranges [18]. Differential voltage analysis (DVA), a similar method, has also been demonstrated to be effective for cell level SOH estimation [19].
Figure 2: Overview of Proposed Module Level State of Health Estimation Algorithm
Battery packs consist of multiple cells in parallel or series for high-energy and high-power applications [20]. In this paper, a battery module refers to a set of battery cells connected in parallel, and the module level SOH is defined as:
@@
For module level SOH estimation, ICA- and DVA-based methods originally developed for the cell level, however, cannot be directly applied, when only module level measurements are available in practice [21]. Inevitable cell-to-cell variations within modules, in terms of cell capacities, internal resistances, contact resistances, temperatures, etc., can cause uneven current distribution among individual cells and, thus, distort the module level IC and DV features [22, 23, 24]. Take IC peak, a feature used for cell level SOH estimation [13, 14, 15, 18], as an example. For lithium nickel-manganese-cobalt oxide 622 (NMC622) modules with three cells connected in parallel, experimental data show that the module level IC peak could be the same for modules with different module level SOH (Fig. 0(a)) or different for modules with the same module level SOH (Fig. 0(b)). Thus, IC peaks alone can no longer be used for module level SOH estimation.
Despite the aforementioned issues, it is a common practice to apply the cell level ICA- and DVA-based methods directly to estimate module level SOH [24, 17], while acknowledging the negative impacts of cell-to-cell variations on the methods. Thus, developing a systematic, robust, and generalizable method for estimating module level SOH under cell-to-cell variations is an open problem.
@@
In this paper, a systematic and generalizable method, as shown in Fig. 2, is proposed to estimate module level SOH with high accuracy and confidence in the presence of cell-to-cell variations. The contributions of the paper are four-fold:
@@
  • First, a new information theory-based feature selection algorithm is proposed to find an optimal set of features for module level SOH estimation. The feature selection process is independent of subsequent learning algorithms, while the selected features reveal physical insights about module level SOH.
@@
  • Second, a new relevance vector regression (RVR)-based SOH estimation model is developed to not only extract point estimates for SOH but also obtain three-sigma credible intervals. The sparsity of the model is automatically imposed, which leads to low model complexity and makes it suitable for onboard implementations.
@@
  • Third, applied to a large dataset of NMC622 modules, the proposed method achieves 0.5% root-mean-square error (RMSE) and 1.5% average three-sigma value for module level SOH estimation, which has significant improvements compared to the results (1.3% RMSE and 3.9% average three-sigma value) obtained by directly applying the conventional IC peak-based method originally developed for the cell level.
@@
  • Fourth, the proposed method has small computational requirements for onboard implementations.
@@
To elucidate the proposed method in Fig. 2, the paper is organized as follows. Section II summarizes all the IC and DV features related to SOH in general. Section III develops the new information theory-based feature selection algorithm. Section IV presents the new RVR-based SOH estimation model. Section V discusses the battery datasets used in this study. Section VI demonstrates how to interpret the proposed method using experimental battery cell data. Section VII assesses the performance of the proposed method using experimental battery module data. The onboard computational requirements involved with the implementation of the proposed method are discussed in Section VIII. Section IX summarizes the paper.
II BATTERY MODULE LEVEL IC AND DV FEATURES
@@
This section deals with Step 2 of Fig. 2 where all the available features in module level IC and DV curves are extracted. Given measured module level charged capacity-voltage ($Q_{c}$-$V$) profiles under constant-current regimes, the module level IC and DV are defined as $\text{IC}=dQ_{c}/dV$ and $\text{DV}=dV/dQ_{c}$, respectively. In this study, support vector regression (SVR) is used to fit the measured profiles and, then, module level IC and DV curves are obtained through analytical derivative calculations to mitigate noise sensitivity caused by numerical derivatives [13, 14, 15].
The following features, as defined graphically in Fig. 3, can be extracted from the IC and DV curves:
  • IC Features: As shown in Fig. 2(a), the $x$- and $y$-coordinates of IC peaks and valleys are their locations and heights, respectively. IC peak areas 1, 2, and 3 are the areas from the minimum voltage to the first IC valley location, from the first to the second IC valley locations, and from the second IC valley location to the maximum voltage, respectively [25]. Furthermore, IC peak partial areas can be defined as the area either above a user-defined horizontal cutoff line or within a user-defined symmetric voltage window around a target IC peak [26], as shown in Fig. 2(b).
  • DV Features: Similarly, DV features, as defined in Fig. 2(c), include the locations ($x$-coordinates) and heights ($y$-coordinates) of DV peaks and valleys.
  • Charging Condition Features: Temperatures and C rates of constant-current charging regimes affect IC and DV curves. Thus, they are also important features for SOH estimation [18].
@@ C Rate [C] -
Table I summarizes all the available features from ICA, DVA, and charging conditions. Their physical meanings and relations to battery degradation are discussed in the references in Table I. Note that, depending on battery chemistries and charging conditions, some features in Table I might not be available in practice, resulting in different numbers of features for different datasets (as to be observed in Section V).
@@ III INFORMATION THEORY-BASED FEATURE SELECTION
The information theory-based feature selection approach is adopted in this work for its computational efficiency [28]. In the context of module level SOH estimation, this approach leverages data to infer underlying physical insights in the presence of cell-to-cell variations. Moreover, its results are independent of subsequent SOH estimation algorithms [28], providing the flexibility for adopting different methods in Step 4 of Fig. 2 for the final SOH estimation.
@@ III-A Definitions
$I\left(F;G\right)=\sum_{f_{i}\in F}\sum_{g_{j}\in G}p\left(f_{i},g_{j}\right)\log{\frac{p\left(f_{i},g_{j}\right)}{p\left(f_{i}\right)p\left(g_{j}\right)}},$ (2)
@@
$I\left(F;G|H\right)$ $=\sum_{h_{q}\in H}\biggl{\{}p\left(h_{q}\right)\cdot$
$\sum_{f_{i}\in F}\sum_{g_{j}\in G}p\left(f_{i},g_{j}|h_{q}\right)\log{\frac{p\left(f_{i},g_{j}|h_{q}\right)}{p\left(f_{i}|h_{q}\right)p\left(g_{j}|h_{q}\right)}}\biggl{\}},$ (3)
where $p(\cdot)$ is the probability distribution function and $i,j,q$ are index symbols used throughout this paper [31]. If $F$, $G$, and $H$ are continuous RVs, summations in these definitions are replaced by integrations [32]. Here, natural logarithms are used. Intuitively, the MI measures the uncertainty reduction of an RV if another RV is known, while the CMI measures the uncertainty reduction of an RV if another RV is known given the third RV [31].
@@ III-B Feature Selection Algorithm
The pseudo-code for the proposed feature selection algorithm is given in Algorithm 1. Symbols used in Algorithm 1 are defined as:
  • $\mathcal{A}$: the set of all features.
  • $\mathcal{S}$: ranked selected feature set
  • $\mathcal{U}$: the set of features not selected yet, $\mathcal{U}=\mathcal{A}\backslash\mathcal{S}$
  • $\mathcal{R}$: removed feature set
$\mathcal{S}$ and $\mathcal{R}$ are the outputs of Algorithm 1.
@@
where ||$\cdot$|| ⋅ | is the cardinality of a set, $X\in\mathcal{U}_{l-1}$ is a feature not selected yet, $X_{j}\in\mathcal{S}_{l-1}$ are selected features, and $Y$ is the output (i.e., SOH). $J(X)$ contains three terms, namely the relevance $\tilde{I}\left(X;Y\right)$, the average redundancy $1/|\mathcal{S}_{l-1}|\cdot\sum_{X_{j}\in\mathcal{S}_{l-1}}\tilde{I}\left(X;X_{j}\right)$, and the average complementarity $1/|\mathcal{S}_{l-1}|\cdot\sum_{X_{j}\in\mathcal{S}_{l-1}}\tilde{I}\left(X;X_{j}|Y\right)$. These terms are explained as follows:
  • Relevance: The relevance of $X$ to $Y$ is defined as the normalized MI between them, i.e., $\tilde{I}\left(X;Y\right)$. The higher the relevance is, the more information is shared between $X$ and $Y$. For cell level SOH estimation, maximizing the relevance picks up features physically related to SOH. For module level SOH estimation, features with high relevance are physically related to SOH and have small sensitivity to cell-to-cell variations.
@@
Optimization (6) is performed until $|\mathcal{U}|=0$.
@@
(On Removing Features) For the SOH estimation problem, the normalized CMI could be high for both complementary and completely redundant features, causing the latter to be selected. Thus, the feature removal step in Algorithm 1 is necessary. Specifically, at one iteration, if $X^{*}$ is selected by solving Optimization (6), then the proposed algorithm removes all the features $X\in\mathcal{U}$ whose redundancies $\tilde{I}\left(X^{*};X\right)$ are above a user-defined threshold $\tilde{I}_{\text{th}}$, i.e., removes the set $\mathcal{D}=\left\{X\in\mathcal{U}:\tilde{I}\left(X^{*};X\right)\geq\tilde{I}_{\text{th}}\right\}$ from $\mathcal{U}$. Thus, these features will not be considered in the next iteration.

### 04 Apr 2024 v4

@@ Abstract
State of health (SOH) estimation for lithium-ion battery modules with cells connected in parallel is a challenging problem, especially with cell-to-cell variations. Incremental capacity analysis (ICA) and differential voltage analysis (DVA) are effective at the cell level, but a generalizable method to extend them to module-level SOH estimation remains missing, when only module-level measurements are available. This paper proposes a new method and demonstrates that, with multiple features systematically selected from the module-level ICA and DVA, the module-level SOH can be estimated with high accuracy and confidence in the presence of cell-to-cell variations. First, an information theory-based feature selection algorithm is proposed to find an optimal set of features for module-level SOH estimation. Second, a relevance vector regression (RVR)-based module-level SOH estimation model is proposed to provide both point estimates and three-sigma credible intervals while maintaining model sparsity. With more selected features incorporated, the proposed method achieves better estimation accuracy and higher confidence at the expense of higher model complexity. When applied to a large experimental dataset, the proposed method and the resulting sparse model lead to module-level SOH estimates with 0.5% root-mean-square errors and 1.5% average three-sigma values. With all the training processes completed offboard, the proposed method has low computational complexity for onboard implementations.
@@
Nomenclature
  • $\mathcal{A}$
    Set of All Features
  • $F,G,H$
    Random Variable
  • $I$
    Mutual Information or Conditional Mutual Information
  • $\tilde{I}$
    Normalized Mutual Information or Normalized Conditional Mutual Information
  • $i,j,l$
    Index
  • $K$
    Kernel Function
  • $N$
    Total Number of Sample Points
  • $Q_{c}$
    Charged Capacity
  • $\mathcal{R}$
    Removed Feature Set
  • $\mathcal{S}$
    Ranked Selected Feature Set
  • $\mathcal{U}$
    Set of Features Not Selected Yet
  • $V$
    Voltage
  • $\boldsymbol{w}$
    Weight Vector
  • $X$
    Feature Random Variable
  • $\boldsymbol{x}$
    Input Vector
  • $Y$
    Output Random Variable
  • $\boldsymbol{y}$
    Vector Containing All Outputs in the Dataset
  • $Z$
    Any Random Variables or Combinations of Random Variables
  • $\boldsymbol{\alpha}$
    Vector with $\alpha_{i}$ where $\alpha_{i}^{-1}=$ Variance of $i$-th Weight Distribution
  • $\beta$
    Scalar where $\beta^{-1}=$ Variance of Noise Distribution
  • $\boldsymbol{\mu}$
    Mean Vector
  • $\boldsymbol{\Phi}$
    Kernel Matrix
  • $\boldsymbol{\Sigma}$
    Covariance Matrix
@@ I INTRODUCTION
Lithium-ion battery state of health (SOH) estimation with onboard measurements is an active research topic for electrified vehicles [1]. SOH, representing the degradation status of batteries, is critical information for battery management systems [2]. Depending on the applications, SOH may be defined as resistance rising, capacity fading, or their combinations [3, 4]. This paper focuses on capacity fading, and degradation in SOH results in the range reduction of electrified vehicles [4].
While there is a rich set of methodologies and results in the literature for SOH estimation, most of the work reported focuses on the cell level [2, 3, 5], using model-based and data-driven methods. Popular model-based methods include Bayesian filters [6, 7, 8], sliding-mode observers [9], and least-square techniques [10]. However, given complicated degradation processes, robust and accurate estimation is difficult to achieve with simple phenomenological or physical models using only onboard measurements. On the other hand, data-driven methods have the advantage of not requiring physics-based models. Various data-driven methods include convolutional and recurrent neural networks [11, 12]. Without insights and structures from physics-based models, however, the effectiveness of these methods solely relies on the comprehensiveness and quality of training data.
@@
Figure 1: Example Module-Level IC Curves around an IC Peak under Cell-to-Cell Variations
One promising middle ground between model-based and data-driven methods is incremental capacity analysis (ICA) [13, 14, 15]. Peaks in incremental capacity (IC) curves reflect degradation mechanisms inside batteries [16, 17]. Thus, ICA incorporates aging physics without using complicated mechanistic models. The capability to provide physical insights about degradation is the key strength of ICA over deep neural network-based methods and other data-based approaches. Furthermore, this method leverages data-driven techniques to correlate physical features (IC peaks) and SOH, and uses the learned model to estimate SOH [18]. ICA has been shown to be effective for different battery chemistries under different temperatures, C rates, and charging ranges [18]. Differential voltage analysis (DVA), a similar method, has also been demonstrated to be effective for cell-level SOH estimation [19].
Figure 2: Overview of Proposed Module-Level State of Health Estimation Method
Battery packs consist of multiple cells in parallel or series for high-energy and high-power applications [20]. In this paper, a battery module refers to a set of battery cells connected in parallel, and the module-level SOH is defined as:
@@
For battery modules with cells connected in parallel, inevitable cell-to-cell variations within modules exist in cell capacities, internal resistances, contact resistances, temperatures, etc. [21, 22]. Different papers in the literature have reported contradictory results about the evolution of cell-to-cell variations within modules over time. Take cell-to-cell variations in capacities and internal resistances as examples. [23] and [24] have reported that capacities and internal resistances of individual cells within modules converge due to self-balancing, whereas [21] and [22] have reported diverging results. Thus, unless the research community comes to a solid conclusion about how cell-to-cell variations within modules will evolve, the module-level SOH estimation must be performed under the assumption that the cell-to-cell variations exist within battery modules.
@@
For module-level SOH estimation, however, cell-to-cell variations within modules hinder ICA- and DVA-based methods [25]. Cell-to-cell variations cause uneven current distribution among individual cells and, thus, distort the module-level IC and DV features [26]. Take IC peak, a feature used for cell-level SOH estimation [13, 14, 15, 18], as an example. For lithium nickel-manganese-cobalt oxide 622 (NMC622) modules with three cells connected in parallel (208Ah nominal module capacity), experimental data show that the module-level IC peak could be the same for modules with different module-level SOH (Fig. 0(a)) or different for modules with the same module-level SOH (Fig. 0(b)) under 86A constant-current charging and room temperature. Thus, the correlation between IC peaks and SOH is obscured by cell-to-cell variations. Some features could be found robust to cell-to-cell variations, but these are only phenomenological observations and the insensitivity of these features could disappear when battery chemistries or charging conditions change. For example, IC peaks and distances between two DV inflection points are found robust to cell-to-cell variations in [26] and [27], respectively, for lithium iron phosphate (LFP) modules under their specific conditions. However, for the NMC622 modules discussed above, the IC peak changes significantly due to cell-to-cell variations, while only one DV inflection point exists at the module level. Thus, a generalizable method to find optimal module-level IC and DV features for module-level SOH estimation is still missing.
@@
In this paper, a new and generalizable method, as shown in Fig. 2, is proposed to estimate module-level SOH with high accuracy and confidence in the presence of cell-to-cell variations. Instead of selecting features based on phenomenological observations as in the existing approaches, this paper proposes an algorithm that can find optimal features for module-level SOH estimation. The optimal features can change when situations change, but the proposed algorithm stays the same. Thus, the proposed algorithm is generalizable. Specifically, the contributions of the paper are four-fold:
@@
  • First, an information theory-based feature selection algorithm is proposed to find an optimal set of features for module-level SOH estimation. The feature selection process is independent of subsequent learning algorithms. To the best of our knowledge, it is the first application of the information theory-based feature selection to module-level SOH estimation.
@@
  • Second, a relevance vector regression (RVR)-based SOH estimation model is developed to not only extract point estimates for SOH but also obtain three-sigma credible intervals. The sparsity of the model is automatically imposed, which leads to low model complexity and makes it suitable for onboard implementations.
@@
  • Third, applied to a large dataset of NMC622 modules, the proposed method achieves 0.5% root-mean-square error (RMSE) and 1.5% average three-sigma value for module-level SOH estimation, which has significant improvements compared to the results (1.3% RMSE and 3.9% average three-sigma value) obtained by directly applying the conventional IC peak-based method originally developed for the cell level.
@@
  • Fourth, the proposed method is demonstrated to have a low computational requirement, making it feasible for onboard implementations.
@@
To elucidate the proposed method in Fig. 2, the paper is organized as follows. Section II summarizes all the IC and DV features related to SOH in general. Section III develops the proposed information theory-based feature selection algorithm. Section IV presents the proposed RVR-based SOH estimation model. Section V discusses the battery datasets used in this study. Section VI demonstrates how to interpret the proposed method using experimental battery cell data. Section VII assesses the performance of the proposed method using experimental battery module data. The onboard computational requirements involved with the implementation of the proposed method are discussed in Section VIII. Section IX summarizes the paper.
II BATTERY MODULE-LEVEL IC AND DV FEATURES
@@
This section deals with Step 2 of Fig. 2 where all the available features in module-level IC and DV curves are extracted. Given measured module-level charged capacity-voltage ($Q_{c}$-$V$) profiles under constant-current regimes, the module-level IC and DV are defined as $\text{IC}=dQ_{c}/dV$ and $\text{DV}=dV/dQ_{c}$, respectively. In this study, support vector regression is used to fit the measured profiles and, then, module-level IC and DV curves are obtained through analytical derivative calculations to mitigate noise sensitivity caused by numerical derivatives [13, 14, 15].
The following features, as defined graphically in Fig. 3, can be extracted from the IC and DV curves. For IC features, as shown in Fig. 2(a), the $x$- and $y$-coordinates of IC peaks and valleys are their locations and heights, respectively. IC peak areas 1, 2, and 3 are the areas from the minimum voltage to the first IC valley location, from the first to the second IC valley locations, and from the second IC valley location to the maximum voltage, respectively [28]. IC peak partial areas can be defined as the area either above a user-defined horizontal cutoff line or within a user-defined symmetric voltage window around a target IC peak [29], as shown in Fig. 2(b). Similarly, DV features, as defined in Fig. 2(c), include the locations ($x$-coordinates) and heights ($y$-coordinates) of DV peaks and valleys. For charging condition features, temperatures and C rates of constant-current charging regimes affect IC and DV curves. Thus, they are also important features for SOH estimation [18].
Table I summarizes all the available features from ICA, DVA, and charging conditions. Their physical meanings and relations to battery degradation are discussed in the references in Table I. Note that, depending on battery chemistries and charging conditions, some features in Table I might not be available in practice, resulting in different numbers of features for different datasets (as to be seen in Section V).
@@ III INFORMATION THEORY-BASED FEATURE SELECTION
The information theory-based feature selection approach is adopted in this work for its computational efficiency [31]. In the context of module-level SOH estimation, this approach leverages data to infer underlying physical insights in the presence of cell-to-cell variations. Moreover, its results are independent of subsequent SOH estimation algorithms [31], providing the flexibility for adopting different methods in Step 4 of Fig. 2 for the final SOH estimation.
@@ III-A Definitions
$I\left(F;G\right)=\sum_{f\in F}\sum_{g\in G}p\left(f,g\right)\log{\frac{p\left(f,g\right)}{p\left(f\right)p\left(g\right)}},$ (2)
@@
$I\left(F;G|H\right)$ $=\sum_{h\in H}\biggl{\{}p\left(h\right)\cdot$
$\sum_{f\in F}\sum_{g\in G}p\left(f,g|h\right)\log{\frac{p\left(f,g|h\right)}{p\left(f|h\right)p\left(g|h\right)}}\biggl{\}},$ (3)
where $p(\cdot)$ is the probability distribution function [34]. If $F$, $G$, and $H$ are continuous RVs, summations in these definitions are replaced by integrations [35]. Here, natural logarithms are used. Intuitively, the MI measures the uncertainty reduction of an RV if another RV is known, while the CMI measures the uncertainty reduction of an RV if another RV is known given the third RV [34].
@@ III-B Feature Selection Algorithm
The pseudo-code for the proposed feature selection algorithm is given in Algorithm 1. Symbols used in Algorithm 1 are defined as: $\mathcal{A}$ – the set of all features, $\mathcal{S}$ – ranked selected feature set, $\mathcal{U}$ – the set of features not selected yet and $\mathcal{U}=\mathcal{A}\backslash\mathcal{S}$, $\mathcal{R}$ – removed feature set. $\mathcal{S}$ and $\mathcal{R}$ are the outputs of Algorithm 1.
@@
where ||$\cdot$|| ⋅ | is the cardinality of a set, $X\in\mathcal{U}_{l-1}$ is a feature not selected yet, $X_{j}\in\mathcal{S}_{l-1}$ are selected features ($j$ is an index), and $Y$ is the output (i.e., SOH). $J(X)$ contains three terms, namely the relevance $\tilde{I}\left(X;Y\right)$, the average redundancy $1/|\mathcal{S}_{l-1}|\cdot\sum_{X_{j}\in\mathcal{S}_{l-1}}\tilde{I}\left(X;X_{j}\right)$, and the average complementarity $1/|\mathcal{S}_{l-1}|\cdot\sum_{X_{j}\in\mathcal{S}_{l-1}}\tilde{I}\left(X;X_{j}|Y\right)$. These terms are explained as follows:
  • Relevance: The relevance of $X$ to $Y$ is defined as the normalized MI between them, i.e., $\tilde{I}\left(X;Y\right)$. The higher the relevance is, the more information is shared between $X$ and $Y$. For cell-level SOH estimation, maximizing the relevance picks up features physically related to SOH. For module-level SOH estimation, features with high relevance are physically related to SOH and have small sensitivity to cell-to-cell variations.
@@
Thus, for module-level SOH estimation, at each iteration, the proposed algorithm selects the feature based on whether it provides the best trade-off among relevance, redundancy, and complementarity, not on how much it is distorted by cell-to-cell variations. Optimization (6) is performed until $|\mathcal{U}|=0$.
@@
(On Removing Features) For SOH estimation problems, normalized CMI could be high for both complementary and completely redundant features, causing the latter to be selected. Thus, the feature removal step in Algorithm 1 is necessary. Specifically, at one iteration, if $X^{*}$ is selected by solving Optimization (6), then the proposed algorithm removes all the features $X\in\mathcal{U}$ whose redundancies $\tilde{I}\left(X^{*};X\right)$ are above a user-defined threshold $\tilde{I}_{\text{th}}$, i.e., removes the set $\mathcal{D}=\left\{X\in\mathcal{U}:\tilde{I}\left(X^{*};X\right)\geq\tilde{I}_{\text{th}}\right\}$ from $\mathcal{U}$. Thus, these features will not be considered in the next iteration.